OpenStudy (inkyvoyd):

Find a sixth-degree Maclaurin polynomial for the function \(g(x)=\frac{1}{(x^2+1)^\frac{5}{2}}\)

OpenStudy (inkyvoyd):

any easy way to do this, or do I just have to grind out 6 derivatives?

OpenStudy (nubeer):

ii dont think there is any easy way.. u have to find six derivatives..

OpenStudy (inkyvoyd):

@nubeer after the second derivative it gets really nasty because the 2x comes out from the chain rule, and then I have to use the product rule...

OpenStudy (nubeer):

well when i use to do those i had to normlaly find up to 4th derivative because teachers new it get nasty as move on to higher derivatives.

OpenStudy (inkyvoyd):

since it's maclaurin series though whenver the term in front has an x in it it becomes 0, which makes things easier but it's still a pain.

OpenStudy (nubeer):

well simplification is easy but i think this kind of question normally would carry marks for showing the derivatives.

OpenStudy (inkyvoyd):

\(g'(0)=(0^2+1)^{-7/2}*\frac{-5}{2}*2*0\)

OpenStudy (inkyvoyd):

Yeah @nubeer these are practice problems for the AP test (free response section) so I am assuming that if I don't do it right I"ll get docked

OpenStudy (nubeer):

lol maybe.. but it will certainly would mkae u good at taking derivatives ;)

OpenStudy (nubeer):

the first term i think is 0 but later on i think when u start taking product rule u migh have terms which wont be turning into 0

OpenStudy (inkyvoyd):

yeah for g''(x) I have a term that isn't 0

OpenStudy (inkyvoyd):

calculating g'''(x) right now >.>

OpenStudy (amistre64):

maybe fine 1/u^(5/2) and subin the innards afterwards

OpenStudy (amistre64):

$235 fine should suffice .. itll crack :/

OpenStudy (inkyvoyd):

btw it's a "no calculator" problem -_-

OpenStudy (amistre64):

thats why i keep a slide rule handy

OpenStudy (inkyvoyd):

LOL what are you talking about I've only heard those in tall tales

OpenStudy (amistre64):

those old dusty journals of mathematics .... i know we are told they are just folk lore and childrens dreams .. but i feel deep down in my gut that there is some truth to them ;)

OpenStudy (amistre64):

maclaurin is f^(n)(0) right?

OpenStudy (inkyvoyd):

yes.

OpenStudy (anonymous):

i got one at a yard sale. looks great but i have no idea how to use it. uses logs somehow

OpenStudy (inkyvoyd):

on the bright side I just realized that u'''(x)=0, and that u'(0), and u''(0)=0. Unfortunately, I am still trying to find g'''(x)

OpenStudy (inkyvoyd):

I see a pattern in the derivative though

OpenStudy (amistre64):

this is just a pain ....

OpenStudy (inkyvoyd):

yeah if I had this question on my AP test I would not be able to finish it in the entire time...

OpenStudy (amistre64):

so far, letting k=-5/2 \[f=1\] \[f'=0\] \[f''=2k\] \[f'''=0\] \[f''''=12k(k-1)\\ \]

OpenStudy (amistre64):

you think f''''' = 0?

OpenStudy (amistre64):

the good news is its fitting so far :) http://www.wolframalpha.com/input/?i=1-5x%5E2%2F2%2B35x%5E4%2F2%2C+y%3D1%2F%28x%5E2%2B1%29%5E%285%2F2%29

OpenStudy (inkyvoyd):

sorry I was in school and had to leave...